COS 36 DIGREES=?
PLEACE TELL THE ANSWER BROTHERS VERY SPEED.

Grade:11

3 Answers

Vijay Mukati
8 years ago
Dear Student,

Assume, x = 18, so 2*x = 36 and 3*x = 54.
Since 2x = 90 – 3x
or sin (2x) = sin(90-3x) = cos 3x
or 2sinx*cosx = 4cos3x – 3cosx,
From here find the value of either cosx or sinx,
Then cos36 = cos 2x = 1 – 2sin2x.

Follow the above steps.

Thanks.
Mounik
8 years ago
Orey VP cos 36= (1+ √5)/4
This is the answer for Cos 36 If u want any answes ask me i am available any time
SAI SARDAR
8 years ago
Dear Sreekanth Indla,
               The value of cos 36 is 0.8090..............................................................................................If you like this please approve it.Thank you.